please solve step by step, and explain please,

  • Vab = 300/0 V

  • ZP = 10/30 ohms

  • fnd Pa and Pb

А Z B bo um 2 PROBLEMA 13.35
Answer #1

10 (30s (1) mm PC Teleser 6 PC Zp for womentes A free = al Upc = vac **THI p PA = Upc. lice. Ces (LV per Icel PA = 7. Jac c- ? كم /N res 300 1 J3 L-30 lo [30 30 J3 L-60° 263 Wlas 300 T 53 L-150 30 13 10 L 30 3 1-10 A ic = 3oo > In _-270 30 ड L-300Po = 76. Voc. Ces og m o 30 L-10c, Voc = 300 E-roo OB 75 ņ -poc RRC 03-60 Voc 0 g -609 PB 30 300 ) CesGo 7598.076 W JB Po =Thank you.